# A lightweight and robust method for blind wideband-to-fullband extension of speech

### v2

@@ Abstract
Reducing the bandwidth of speech is common practice in resource constrained environments like low-bandwidth speech transmission or low-complexity vocoding. We propose a lightweight and robust method for extending the bandwidth of wideband speech signals that is inspired by classical methods developed in the speech coding context. The resulting model has just $\sim 370$ K parameters and a complexity of $\sim 140$ MFLOPS (or $\sim 70$ MMACS). With a frame size of 10 ms and a lookahead of just 0.27 ms the model is well-suited for common wideband speech codecs. We evaluate the model’s robustness by pairing it with the Opus SILK speech codec (1.5 release) and verify in a P.808 DCR listening test that it significantly improves quality from 6 to 12 kb/s. We also demonstrate that Opus 1.5 together with the proposed bandwidth extension at 9 kb/s meets the quality of 3GPP EVS at 9.6 kb/s and that of Opus 1.4 at 18 kb/s showing that the blind bandwidth extension can meet the quality of classical guided bandwidth extensions.
I Introduction
@@
While bandwidth reduction is effective for saving resources and (mostly) maintains speech intelligibility, it also degrades the listening experience and can lead to listener fatigue. Therefore, a blind bandwith extension (BWE) method can have a largely positive impact for billions of listeners every day. However, low complexity is critical for the applications stated above as typical target devices like smartphones can have rather limited compute. Furthermore, robustness is essential since any real-world deployment faces huge variability of input signals.
@@
BWE is a well-studied topic and both classical ([18, 8, 29, 28, 4]) and DNN-based ([10, 23, 17, 19, 13]) methods have been proposed for this task. While classical methods are very low in complexity, they struggle with blind highband estimation and are therefore most effective when provided with a bit of side information. DNN-based methods, on the other hand, are much better at high-band modeling, but even dedicated low-complexity algorithms still operate in the range of multiple GFLOPS (e.g. $\sim 13$ GFLOPS in [23] or $\sim 7$ GFLOPS in [13]) which prevents their deployment on smaller devices.
@@
In this paper we seek to overcome this problem by combining the high-band modeling capacity of data-driven, DNN-based methods with the simplicity and low complexity of DSP-based BWE methods. The approach is inspired by classical time-domain bandwidth extension, where a bandwidth-extending operation like non-linear function application or spectral folding is applied to the upsampled signal and combined with time-varying spectral shaping filters. The signal-processing part of the resulting algorithm only consists of classical DSP, i.e. fixed non-linear mapping, fixed and time-varying linear filtering and time-varying sample-wise weighting. The time-varying filters and sample-wise weights in turn are produced by a small DNN which thus governs the content and shape of the generated highband signal. The resulting model has $\sim 370$ K parameters and a computational complexity of $\sim 140$ MFLOPS (or $\sim 70$ MMACS) which makes it suitable for use even on older smartphone devices. Furthermore, since it is built around a low-delay upsampler, the signal path induces a delay of only 13 samples at 48 kHz or $\sim 0.27$ ms and the feature path has a framing delay of $10$ ms. This means that if combined with a speech production system that operates on multiples of 10 ms (which most speech codecs do) the total added delay is as low as $\sim 0.27$ ms
@@
To test model robustness, we combine it with the Opus codec (1.5 release) and confirm in a P.808 listening test [^1] The page https://janpbuethe.github.io/BWEDemo contains some demo samples. It also includes vocoding examples which were not included in the listening test. that BWE provides consistent improvement for all tested bitrates. We furthermore include the superwideband codec EVS [3] at 9.6 kbs and Opus 1.4 at 18 kb/s which produces fullband speech in a hybrid coding mode. The test results show that both are statistically tied with the bandwidth-extended Opus 1.5 at 9 kb/s showing that Opus 1.5 with blind bandwidth extension can meet the quality of codecs with classical parametric resp. semi-parametric bandwidth extensions.
Finally, it should be pointed out that the implementation of the proposed BWE method is based on the adaptive DDSP modules AdaConv and AdaShape proposed in [7] and [6]. The successful application to low-complexity BWE demonstrates the general usefulness of this adaptive signal processing approach and there are likely more applications to which it could be applied.
@@
A python implementation is available at https://gitlab.xiph.org/xiph/opus/-/tree/exp_bwe/dnn/torch/osce (BBWENet).
II Model Description
@@
For upsampling, the model leverages the libopus 16 to 48 kHz upsampler [^2] https://gitlab.xiph.org/xiph/opus which operates in two stages: in a first stage the signal is upsampled by a factor two using IIR filters and in a second stage a 1.5x interpolation with short FIR filters is performed. The upsampler is both low in complexity and has a low delay of 13 samples at 48 kHz, which is also the total delay of the signal path on the right hand side of Figure 1. IIR filters are approximated by long FIR filters for training.
@@
For the actual bandwidth extension, the model uses a hybrid approach that utilizes two common methods for time-domain bandwidth extension. The first one is non-linear function application, which has the advantage of generating a consistent harmonic extension if the baseband is quasi-periodic (voiced speech). We deviate from the usual choice of non-linearity (absolute value or rectified linear unit) and choose instead a non-linearity that extends the signal more agressively (1). It is designed to approximately preserve the scale of the signal and to induce a similar amount of distortion regardless of the scale of the input signal.
@@
The second extension method is motivated by spectral folding as proposed in [18]. However, we use the term spectral folding in a broader sense of multiplying the signal with a locally periodic sequence of non-negative weights. We imlement this using the AdaShape module proposed in [6] which multiplies the input signal with a sequence of weights calculated from the sequence of latent feature vectors.
@@
The features for the feature encoder on the left side of Figure 1 are designed to provide the model with basic signal properties like spectral envelope, and pitch and voicing while being simple to compute. We chose to use a Hanning-window STFT with a window size of 20 ms and a hop size of 10 ms from which we compute a 32-band ERB-scale log magnitude spectrogram and instantaneous frequency information for the first 40 STFT bins in the form
$\operatorname{IF}(k,n)=\frac{X(n,k)\,X^{*}(n-1,k)}{\mid X(n,k)\,X^{*}(n-1,k)\mid}.$ (3)
Here, $n$ denotes the frame index, $k$ denotes the frequency index, and denotes complex conjugation. Instantaneous frequencies are included as features since it has been demonstrated that they are sufficient for high-accuracy pitch estimation [24]. The feature encoder upsamples the 72-dimensional feature vector from 100 to 200 Hz and includes a GRU for accumulating context.
III Training
III-A Strategy and Data
Bandwidth extension, viewed as an inverse problem, is generally ill posed since even for the same source signal the actual recording will depend on environmental factors such as acoustic environment or recording device. It is possible to a certain extent to fit a model to the characteristics of a specific homogeneous dataset, but we observed that this can lead to poor generalisation. This is in line with the findings in [15] which in particular highlights the critical impact of microphone channels. The training procedure for the proposed model therefore prioritizes plausibility and robustness over over correctness.
@@ III-A Strategy and Data
To achieve robustness, we train on a mixture of multiple high-quality TTS datasets [9, 16, 22, 11, 14, 20, 26, 12, 5] containing more than 900 speakers in 34 languages and dialects. Since some of the datasets contained upsampled wideband and superwideband recordings, we filter out items with very little energy high frequency range. Furthermore, we apply the following data augmentation steps to increase model robustness to unseen speech:
@@
Step 1) prevents the model from relying too much on low frequencies, which should be irrelevant for extending the wideband signal. Step 2) teaches the model not to extend the noise but only the speech from the baseband. This is an intentional design decision and it stems from the observation, that adding a bandwidth extension to signal of poor quality can reduce perceived quality as the resulting signal sounds noisier than the bandlimited input signal. Steps 2) and 3), when both carried out on the same item, are executed in random order.
@@ III-B Losses and Training
IV Evaluation
IV-A Subjective Evaluation
@@
To test model performance and robustness, we carried out a multi-bitrate speech coding test using the P.808 DCR methodology with samples from the EARS dataset [21] from which no portion was included in the training data. To this end, we extracted three sentence pairs per speaker from the regular speech category. We tested the proposed bandwidth extension both for Opus 1.5 with decoder complexity 10 (which means NoLACE enhancement will be applied) and for clean speech input and report significant improvement for all conditions. In particular, the improvement of Opus 1.5 at 6 kb/s is remarkable since the baseband already exhibits very audible distortions.
As comparison points we included the 3GPP EVS codec at 9.6 kb/s and the results show that Opus with lowband enhancement and blind bandwidth extension matches the quality even at the same bitrate. This indirectly gives a comparison of the proposed blind BWE method to a guided, parametric BWE. Furthermore, we included EnCodec at 6 and 12 kb/s to test the hybrid approach against the full end-to-end neural coding approach. While EnCodec delivers better quality at 6 kb/s, even Opus 1.5 wideband at 9 kb/s already significantly outperforms EnCodec at 12 kb/s. This could be in part a domain shift problem as the originally provided samples at 12 kb/s sound better than the coded samples from the EARS dataset. In particular, a pre-test also included the neural codec AudioDec, which exhibited severe quality degradation on this dataset and was therefore excluded from the final test (examples are included in the demo page). These results suggest, that the hybrid approach may be more robust than the end-to-end approach and they certainly suggest that robustness must be a concern when considering to deploy an end-to-end codec.
@@
Finally, we added Opus 1.4 at 18 kb/s as a self comparison point to evaluate the effect of combined lowband enhancement and highband extension. Opus 1.5 at 9 kb/s is already statistically tied to Opus 1.4 at 18 kb/s and equal quality is likely achieved around 10 kb/s. So in combination, these backward-compatible enhancement methods yield a bitrate reduction of 45 to 50%. Furthermore, this comparison gives a second, indirect comparison of the blind BWE to a Opus-coded highband in hybrid mode.
@@
IV-B Model Inspection
Due to the simple architecture of the signal-processing part of the model, it is straight forward to inspect the contribution of the individual modules to the final bandwidth extension. In particular, since the mixing or post-filtering layers are linear, the second bypass signal $y_{32}(t)$ and the output signal $y_{48}(t)$ can be decomposed as a sum of signals stemming from the previous bypass channel and the output channels of the AdaShape and NonLin modules.
Spectrograms of these contributions are displayed in Figure 3. In the first stage, unvoiced signal parts are primarily extended by the AdaShape module while voiced signal parts are extended by the non-linearity. In the second stage, the SWB to FB extension is primarily constructed from the AdaShape output and from imaging remaining from the short FIR interpolation filters.
@@ IV-B Model Inspection
V Conclusion
We proposed a lightweight method for wideband-to-fullband extension of speech. We demonstrated the method’s effectiveness and robustness by applying it to both coded and clean speech. While the underlying model was only trained on clean speech, the bandwidth extension resulted in significant quality improvement even when paired with a low-bitrate speech codec.

### v3

@@ Abstract
Reducing the bandwidth of speech is common practice in resource constrained environments like low-bandwidth speech transmission or low-complexity vocoding. We propose a lightweight and robust method for extending the bandwidth of wideband speech signals that is inspired by classical methods developed in the speech coding context. The resulting model has just $\sim$370 K parameters and a complexity of $\sim$140 MFLOPS (or $\sim$70 MMACS). With a frame size of 10 ms and a lookahead of only 0.27 ms, the model is well-suited for use with common wideband speech codecs. We evaluate the model’s robustness by pairing it with the Opus SILK speech codec (1.5 release) and verify in a P.808 DCR listening test that it significantly improves quality from 6 to 12 kb/s. We also demonstrate that Opus 1.5 together with the proposed bandwidth extension at 9 kb/s meets the quality of 3GPP EVS at 9.6 kb/s and that of Opus 1.4 at 18 kb/s showing that the blind bandwidth extension can meet the quality of classical guided bandwidth extensions thus providing a way for backward-compatible quality improvement.
1 Introduction
@@
While bandwidth reduction is effective for saving resources and (mostly) maintains speech intelligibility, it does degrade the listening experience and can lead to listener fatigue. Therefore, a blind bandwith extension (BWE) method can have a largely positive impact for billions of listeners every day. However, low complexity is critical for the applications stated above as typical target devices like smartphones can have rather limited compute. Furthermore, robustness is essential since any real-world deployment faces huge variability of input speech signals.
@@
BWE is a well-studied topic and both classical ([5, 6, 7, 8, 9]) and DNN-based ([10, 11, 12, 13, 14]) methods have been proposed for this task. While classical methods often have low complexity, they struggle with blind highband estimation and are therefore most effective when provided with side information. DNN-based methods, on the other hand, are much better at high-band modeling, but even dedicated low-complexity algorithms still operate in the range of multiple GFLOPS (e.g. $\sim$ 13 GFLOPS in [11] or $\sim$7 GFLOPS in [14]) which prevents their deployment on mobile devices.
@@
In this paper we seek to overcome this problem by combining the high-band modeling capacity of data-driven, DNN-based methods with the simplicity and low complexity of DSP-based BWE methods. The approach is inspired by classical time-domain bandwidth extension, where a bandwidth-extending operation like non-linear function application or spectral folding is applied to the upsampled signal and combined with time-varying spectral shaping filters. The signal-processing part of the resulting algorithm only consists of classical DSP, i.e. fixed non-linear mapping, fixed and time-varying linear filtering and time-varying sample-wise weighting. The time-varying filters and sample-wise weights in turn are adapted by a small DNN which governs the content and shape of the generated highband signal. The resulting model is trained with a mixture of regression and adversarial losses. It has $\sim$370 K parameters and a computational complexity of $\sim$140 MFLOPS (or $\sim$70 MMACS) which make it suitable for use even on older smartphone devices. Furthermore, since it is built around a low-delay upsampler, it requires only a lookahead of 0.27 ms in addition to a framing delay of 10 ms that may be shared with the wideband speech production system.
@@
To test model robustness, we combine it with the Opus codec (1.5 release) and confirm in a P.808 listening test [^1] The page https://janpbuethe.github.io/BWEDemo contains some demo samples. It also includes vocoding examples which were not included in the listening test. that the BWE model, though only trained on clean speech, provides consistent improvement for all tested bitrates. We furthermore include the superwideband codec EVS [15] at 9.6 kb/s and Opus 1.4 at 18 kb/s which produces fullband speech in a hybrid coding mode. The test results show that both are statistically tied with the bandwidth-extended Opus 1.5 at 9 kb/s showing that Opus 1.5 with blind bandwidth extension can meet the quality of codecs with classical parametric resp. semi-parametric bandwidth extensions.
@@
A python and C implementation is available at https://gitlab.xiph.org/xiph/opus/-/tree/waspaa_2025_bwe (BBWENet).
2 Model Description
@@
For upsampling, the model leverages the libopus 16-to-48-kHz upsampler [^2] https://gitlab.xiph.org/xiph/opus which operates in two stages: in a first stage the signal is upsampled by a factor two using IIR filters and in a second stage a 1.5x interpolation with short FIR filters is performed. The upsampler has low complexity and induces a low delay of 13 samples at 48 kHz which is also the total delay of the signal path on the right hand side of Figure 1. IIR filters are approximated by sufficiently long FIR filters for training.
@@
For the actual bandwidth extension, the model uses a hybrid approach that utilizes two common methods for time-domain bandwidth extension. The first one is the application of a non-linear function, a method that generates a consistent harmonic extension for quasi-periodic lowbands (voiced speech). We deviate from the usual choice of non-linearity (absolute value or rectified linear unit) and use a non-linearity instead that extends the signal more agressively (1). It is designed to approximately preserve the scale of the signal and to induce a similar amount of distortion regardless of the scale of the input signal.
@@
The second extension method is motivated by spectral folding as proposed in [5]. We use the term spectral folding in a broader sense of multiplying the signal with a locally periodic sequence of non-negative weights which we implement using the AdaShape module proposed in [17]. It multiplies the input signal with a sequence of weights calculated from a sequence of latent feature vectors.
@@
The input features on the left side of Figure 1 are designed to provide the model with basic signal properties like spectral envelope, pitch and voicing while being simple to compute. We chose to use a Hanning-window STFT with a window size of 20 ms and a hop size of 10 ms from which we compute a 32-band ERB-scale log magnitude spectrogram and complex phase difference information for the first 40 STFT bins in the form
$\Delta\Phi(k,n)=\frac{X(n,k)\,X^{*}(n-1,k)}{\mid X(n,k)\,X^{*}(n-1,k)\mid},$ (3)
where $n$ denotes the frame index, $k$ denotes the frequency index, and denotes complex conjugation. The complex phase differences are included as features since they proved sufficient for high-accuracy pitch estimation [18]. The feature encoder upsamples the 72-dimensional feature vector from 100 to 200 Hz and includes a GRU for accumulating context.
3 Training
3.1 Strategy and Data
Bandwidth extension, viewed as an inverse problem, is generally ill posed since even for the same source signal, the actual recording will depend on environmental factors such as acoustic environment or recording device. It is possible to a certain extent to fit a model to the characteristics of a specific homogeneous dataset, but we observed that this can lead to poor generalisation. This is in line with the findings in [19] which in particular highlights the critical impact of microphone channels. The training procedure for the proposed model therefore prioritizes plausibility and robustness over correctness.
@@ 3.1 Strategy and Data
We train the model on a mixture of multiple high-quality TTS datasets [20, 21, 22, 23, 24, 25, 26, 27, 28] containing more than 900 speakers in 34 languages and dialects. Since some of the datasets contained upsampled wideband and superwideband recordings, we filter out items with very little energy high frequency range. Furthermore, we apply the following data augmentation procedures to increase model robustness to unseen conditions:
@@
Augmentation 1) prevents the model from relying too much on low frequencies, which should be irrelevant for extending the wideband signal. Augmentation 2) teaches the model not to extend noise but only the speech from the baseband. This is an intentional design decision and it stems from the observation, that adding a bandwidth extension to a signal of poor quality can reduce perceived quality as the resulting signal sounds noisier than the bandlimited input signal. Augmentations 2) and 3), when both carried out on the same item, are applied in random order.
@@ 3.2 Losses and Training
4 Evaluation
4.1 Subjective Evaluation
@@
To test model performance and robustness, we carried out a multi-bitrate speech coding test using the open source implementation [29] of the P.808 DCR methodology. We sampled three random sentence pairs per speaker from the EARS dataset [30] (regular speech category) as test material and applied loudness normalization. Note that no condition was trained on the EARS dataset. We tested the proposed bandwidth extension both for Opus 1.5 with decoder complexity 10 (which means NoLACE-enhancement [17] will be applied) and for clean speech input and report significant improvement for all conditions. In particular, the improvement of Opus 1.5 at 6 kb/s is remarkable since the baseband already exhibits very audible distortion.
As comparison points we included the 3GPP EVS codec at 9.6 kb/s [^4] EVS codes frequencies up to $14.4$ kHz at this bitrate and the results show that Opus with lowband enhancement and blind bandwidth extension matches the quality even at the same bitrate. This indirectly gives a comparison of the proposed blind BWE method to a guided, parametric BWE. Furthermore, we included EnCodec at 6 and 12 kb/s to test the hybrid approach (mix of DNN and DSP) against the full end-to-end neural coding approach. While EnCodec delivers better quality at 6 kb/s, even Opus 1.5 wideband at 9 kb/s already significantly outperforms EnCodec at 12 kb/s. This likely highlights a lack of robustness to out-of-domain data, as the quality reported in [31] is much higher. In particular, a pre-test also included the neural codec AudioDec, which exhibited severe quality degradation on this dataset and was therefore excluded from the final test (examples are included in the demo page). These results suggest, that the hybrid approach may be more robust than the end-to-end approach and they certainly suggest that robustness must be carefully evaluated when considering to deploy an end-to-end codec.
Finally, we included Opus 1.4 at 18 kb/s as a reference point to evaluate the combined effect of lowband enhancement and highband extension. Opus 1.5 at 9 kb/s is statistically tied to Opus 1.4 at 18 kb/s, with equivalent quality likely achieved around 10 kb/s, which results in a bitrate reduction of 45 to 50%. Additionally, this comparison provides an indirect evaluation of the blind BWE against an Opus-coded highband in hybrid mode.
@@
4.2 Model Inspection
Due to the simple architecture of the signal-processing part of the model, it is straightforward to inspect the contribution of the individual modules to the final bandwidth extension. In particular, since the mixing of the extension layers is linear, the second bypass signal $y_{32}(t)$ and the output signal $y_{48}(t)$ can be decomposed as a sum of signals stemming from the previous bypass channel and the output channels of the AdaShape and NonLin modules.
Spectrograms of these contributions are displayed in Figure 3. It shows that in the first stage, unvoiced signal parts are primarily extended by the AdaShape module while voiced signal parts are mostly extended by the non-linearity. In the second stage, the SWB to FB extension is primarily constructed from the AdaShape output and from imaging remaining from the short FIR interpolation filters.
@@ 4.2 Model Inspection
5 Conclusion
We proposed a lightweight method for wideband-to-fullband extension of speech and demonstrated that it improves wideband speech from multiple sources. In particular, the proposed model enables fullband speech coding with the Opus codec down to very low bitrates without adding additional delay. Our analysis furthermore shows, that the model adopts a dual approach for extending speech, using folding for unvoiced and non-linear extension for voiced speech parts. This insight can potentially help improve classical BWEs in the future. A next step will be to adopt the model to also extend music where first tests showed promising results.